OpenStudy (anonymous):

Multiply and simplify by factoring then write in radical form. 3rd root of (y)^10 multiplied by 3rd root of (16y)^11

OpenStudy (anonymous):

\[\sqrt[3]{y^{10}}*\sqrt[3]{2^{44}y^{11}}=\sqrt[3]{(y^3)^3y}*\sqrt[3]{(2^{14})^3(y^3)^3y^2}\]apply third root all perfect cubes\[=y^3\sqrt[3]{y}*2^{14}y^3\sqrt[3]{y^2}\]commute and associate factors\[=16384y^6\sqrt[3]{y*y^2}=16384y^6\sqrt[3]{y^3}\]\[=16384y^6*y=16384y^7\]

OpenStudy (anonymous):

little correction 14*3=42 16384*y^7*2^(2/3)

OpenStudy (anonymous):

Marina, Would you please write my original problem as you understood it because someone told me I had been writing the problems incorrectly and I want to make sure I wrote this problem correctly.

OpenStudy (anonymous):

thanks marina, so radical form is \[1634y^7\sqrt[3]{4}\](it's the whole arithmetic thing!)

OpenStudy (anonymous):

1. You are welcome. 2. I told you that. 3. You did pretty good job here. 4. It's possible to write in exponent form y^(10/3)*(16y)^(11/3)

OpenStudy (anonymous):

To Mandolino: yes I pretty often check answers there, just to be sure http://www.wolframalpha.com/input/?i=%2816y%29%5E%2811%2F3%29*y%5E%2810%2F3%29

OpenStudy (anonymous):

OO. I didn't pay attention that answer need to be in radical form

OpenStudy (anonymous):

Thanks you guys, but I think you solved the wrong problem. The second radical is not correct. Check my original problem.

OpenStudy (anonymous):

be specific

OpenStudy (anonymous):

the 1634 should be 1638 according to wolfram

OpenStudy (anonymous):

is this your original?\[\sqrt[3]{y^{10}}*\sqrt[3]{(16y)^{11}}\]if so, the answer seems correct with the 1634 becoming 1638

OpenStudy (anonymous):

Yes (*.*), but no bracket around the 16y, but otherwise yes

OpenStudy (anonymous):

i'll go to the current post since it is shorter and will respond better